OpenStudy (anonymous):

focus(-4,-4); opens left; contains (-7,0) write the equation for the parabola given set characteristic

OpenStudy (anonymous):

Any idea?

OpenStudy (anonymous):

The standard form of a left/right open parabola is \((y -k)^2=4p(x-h)\) where vertex (h, k) and focus (h+p, k)

OpenStudy (anonymous):

You have focus (-4,-4) it means h+p= -4 and k =-4 Now, the parabola passes through the point (-7,0) , just replace x=-7 , y =0, k =-4 into the standard form, you have \((0-(-4))^2=4p(-7-h)\) accompany with \(h+p =-4\)

OpenStudy (anonymous):

You have 2 equations, 2 unknowns h, p--> solveable, hence: solve for h, p, then plug them back to standard form of parabola and you are done